### v1

@@ 3.1 Matrix algebra for datalog
We introduce datalog programs and concepts relating to our BMLP framework based on LP_book ; ILP_foundation . A variable is a character string beginning with an uppercase letter. A function, predicate or constant symbol is a character string starting with a lowercase letter. A term is either a variable, a constant symbol, or a function symbol attached to a tuple of terms. A ground term contains no variables. An atom is a predicate symbol applied to a tuple of terms e.g. $r(a,Y)$ where $r$ is a predicate symbol, $a$ is a constant symbol and $Y$ is a variable. An atom is ground when it contains only ground terms. A Horn clause is a disjunctive clause with at most one positive literal. A definite clause is a Horn clause with exactly one positive literal and a definite program is a set of definite clauses. A query is a user-posted Horn clause without any positive literal. A general logic program engine such as a Prolog interpreter can evaluate a query by carrying out a refutation proof of its negation via SLD-resolution (Selective Linear Definite clause resolution). A unit ground clause (fact) is a definite clause with no body literal and no variables. Datalog programs are definite programs without function symbols. A first-order definite clause has the form, $A_{0}\leftarrow A_{1},...,A_{n}$ where $A_{i}=r_{i}(x_{1},...,x_{k})$, $r_{i}$ is a predicate symbol, $x_{j}$ is a term. The arity of a function or predicate symbol is its number of arguments, e.g. $r_{i}/k$ indicates $r_{i}$ has arity $k$. The literal $A_{0}$ is called the head of the clause and the literals $A_{i}$ are the body.
@@
A recursive program has a body literal that appears in the head of a clause. Specifically, we focus on linear and immediately recursive programs as simple recursive programs. A linear and immediately recursive program is a single recursive clause with the recursive predicate appearing in the body only once ioannidis_computation_1986 . This subset of recursive logic programs can represent the relational structure of reaction pathways. This is further elaborated in Section 4. An example of a linear and immediately recursive program involving one non-recursive predicate is:
@@
where unit ground clauses $r_{1}$ and a recursive clause $r_{2}$ can be represented as matrices R1, R2 such that $(\textbf{R}_{k})_{i,j}$ = 1 if $\mathcal{P}\models r_{k}(c_{i},c_{j})$ for constant symbols $c_{i},c_{j}$ and $(\textbf{R}_{k})_{i,j}$ = 0 otherwise. Addition “+” between two boolean matrices is defined as (A + B)i,j = Ai,j $\lor$ Bi,j. Multiplication “$\times$” is defined as $(\textbf{A}\times\textbf{B})_{i,j}$ = $\bigvee_{k=1}^{n}$ Ai,k $\land$ Bk,j and is abbreviated as (AB).
@@ 3.2 Elementary nets
A marking $m^{\prime}$ is reachable from a marking $m$ if and only a finite step sequence exists. A place is reachable from $m$ if and only if $m^{\prime}$ is a reachable marking from $m$ and $m^{\prime}(p)=1$. A marking can be written as a row vector ${[}1,0,1,...{]}$ where each element indicates if the respective place has a token. A transition can also be represented by a row vector ${[}z_{1},z_{2},...,z_{k}{]}$ where $z_{i}$ reflects the assignment of tokens at the pre-set and post-set places. The consecutive firings of transition update markings and the global state of the system. This represents the dynamic changes in the distribution of resources.
@@ 4.1 Modelling a GEM with an OEN
We discuss the relationship between GEMs and OENs. Petri nets are commonly used formalism for representing complex biological systems such as genome-scale metabolic network models and diagnosing their dynamic behaviours sahu_advances_2021 . A genome-scale metabolic network model contains biochemical reactions of substrates. Reactants are substrates initially involved in a reaction and the products are the chemicals yielded by a reaction. A reaction is reversible when the conversion of reactants to products and products to reactants can take place simultaneously. Reversible reactions are commonly treated as two reactions in opposite directions. A reaction involves chemical substrates $x_{i}$ and $y_{j}$:
@@
Given an auxotrophic mutant experiment, we model a GEM as an OEN to find the set of synthesisable substrates by following the metabolic reaction pathways. This requires computing the necessary set of places for transitions to fire. Transitions need to be enumerated to examine if their necessary conditions can ever be met for them to fire. Every simulation of GEM involves a definition of the growth medium consisting of essential chemicals for cellular metabolism such as amino acids, inorganic salts and carbon sources. The definition of a growth medium serves as the basis of metabolic processes and gives an initial marking of OEN. Reachable places extending this marking are additional synthesisable substrates.
@@ 4.2 Transforming an OEN into datalog
Each auxotrophic mutant experiment is treated as a datalog query to find all reachable places based on an initial marking. We evaluate queries of OEN reachability by transforming the OEN into a linear and immediately recursive datalog program. A datalog program consisting of clauses with arity two can be considered a directed graph where edges joining nodes represent the relation between terms. However, the multiplicity of the involved entities in a relation cannot be expressed this way. Instead, places and transitions are represented as constant and predicate symbols. Each argument of our transformed datalog program is a concatenation of constant symbols, e.g. $h2\_o2$. For our specific purpose, since all place nodes in the OEN are related to metabolites, every new concatenated constant symbol describes a set of metabolic substrates.
This transformation with constant symbolic concatenation generalises from a directed graph into a hypergraph berge_hypergraphs_1989 , a directed graph where an edge can connect to any number of nodes. Formally, a hypergraph $G=(V,E)$ is a tuple containing a set of nodes $V$ and a set of edges $E\subseteq 2^{V}\times 2^{V}$ between arbitrary subsets of $V$. Hypernodes are sets of nodes connected by such edges. The incidence graph of a hypergraph is bipartite in that for $v\in V$ and $e\in E$, $e$ is connected to $v$ if and only if $v$ and $e$ are nodes connected by an edge in the bipartite graph. Given an OEN, we consider a hypergraph with an incidence graph of the same bipartite structure. A hypernode represents a set of place nodes and is created for every transition’s input places from the OEN. Hypernodes are denoted by the new constant symbols that concatenate the member places’ names, e.g. a hypernode $\{H_{2},O_{2}\}$ is represented by the constant symbol $h2\_o2$. The number of hypernodes we create is fixed and limited by the number of transitions. This is because each reaction has a finite number of reactants and products. The dynamic behaviour of an OEN is specified via an initial marking. This mapping between an OEN and a hypergraph allows us to write a datalog program. A recursive clause is added which describes a path in the hypergraph.
@@
The right-hand-side figures are the OEN from from Figure 2 and the hypergraph represented by $\mathcal{P}_{1}$. Both graphs show that the place $c_{2}$ is immediately reachable from the initial marking $m_{1}=\{c_{1},c_{2}\}$. The hypernode $\{c_{3},c_{4}\}$ is reachable from via the transition $t_{1}$. The place $c_{5}$ is reachable in two steps via transitions $t_{1}$ and $t_{2}$. We can compute these by evaluating groundings of $pathway(m1,Y)$ in the Herbrand model $M(\mathcal{P}_{1})$ which gives us $pathway(m1,c2)$, $pathway(m1,c3\_c4)$ and $pathway(m1,c5)$.
@@ 5.1.1 Creating boolean matrices
We show boolean matrices created from $\mathcal{P}_{1}$ in Example 1. Individual substrates are represented by constant symbols $c1$, $c2$ and $c3$. Sets of chemical substrates are denoted by additional concatenated constant symbols, $c1\_c2$, $c3\_c4$. On the right-hand side, we create two boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ from the reaction clauses in $\mathcal{P}_{1}$. Since the first reaction_1 clause also contains the constant symbols of the other reaction_1 clauses, to avoid redundancy, only one binary code is created for these clauses in $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
@@
We show the mapping of a single row matrix from the token definition from $\mathcal{P}_{1}$ in Example 1. Individual substrates are represented by constant symbols $c1$ and $c2$. The set of chemical substrates is denoted by an additional concatenated constant symbol, $c1\_c2$.
@@ 5.2 Iterative extension (BMLP-IE)
Theorem 5.1 shows that BMLP-IE has a polynomial runtime cubic in the size of the boolean matrices. Evaluating the program $\mathcal{P}$ using SLD-resolutions has the same time complexity. For $\mathcal{P}$, the size of the subset of the Herbrand base provable from a query $r(u,Y)$ is bounded by $O(n)$. Each member of this subset can be found via a depth-first search. Since there are $k$ unit ground clauses and each variable can be grounded by at most $n$ constants, this query requires at most $n$ search in a tree of $O(k\times n)$ nodes. Therefore, the runtime of BMLP-IE primarily benefits from bit-wise operations. For OEN reachability queries, it is important to note that the size of boolean matrices grows linearly with the number of hypernodes because they are both bounded by the number of transitions.
@@ 6.2 Expected cost of experiments
Experiment planning is often restricted by time and resource costs. A rule of thumb for experiment design is to achieve optimal empirical outcomes within a finite budget. The problem of designing experiments to support or refute scientific hypotheses involves a series of binary decisions. This process is analogous to playing the game of “Guess Who” where players ask a sequence of “yes” or “no” questions whose answers binarily partition the candidate hypotheses. The outcome of an experiment $t$ splits hypotheses $H$ into consistent hypotheses $H_{t}$ and inconsistent hypotheses $\overline{H_{t}}$. The path of selection and outcomes of experiments can be represented as a binary decision tree. To construct an optimal tree, an active learner should seek to label instances consistent with up to half of the hypotheses mitchell_generalization_1982 .
@@ 6.3 Active learning sample complexity
The minimal reduction ratio of an actively selected instance can be computed before it is labelled. The optional selection strategy is to select instances with minimal reduction ratios close to $\frac{1}{2}$ or 50%. In reality, instances might not be as discriminative. To capture the sample complexity advantage of active learning over random example selection, we describe the following bound of an active version space learner’s sample complexity. We consider a learner using random example selection a passive learner since it does not have control over the training examples it uses.
@@ 7.1 Encoding a metabolic network
Our encoding of a metabolic network is based on the Robot Scientist King04:RobotScientist ; ase_progol . The Robot Scientist is implemented using the ILP system Progol progol and finds the synthesisable set of metabolites by traversing a metabolic network. It uses a metabolic network as a deductive database and answers a query of user-defined gene deletion and experimental conditions based on SLD-resolutions. Its deductive database is written in Prolog syntax and uses Prolog list operations to manipulate constant symbols representing metabolic substrates. Metabolic reactions are enumerated based on the set of substrates currently synthesisable. The product substrates of a possible metabolic reaction extend the list of synthesisable substrates. Since the number of metabolic reactions is finite, the closure of members in this list needs to contain essential substrates for the growth of an organism. In an auxotrophic mutant experiment outcome prediction, a gene function is removed. Some key metabolic reaction pathways might no longer be viable due to the removal, leading to an insufficient set of substrates for growth. In this case, the predicted experiment’s outcome is a lack of growth and a phenotypic effect.
In the Robot Scientist, the metabolic network is defined by $enzyme/6$ clauses where each contains an artificial identifier for an enzyme, gene locus identifiers, reaction identifiers, a reaction reversibility indicator and reactant and product identifiers. A gene locus corresponds to a physical site within a genome and is uniquely associated with a gene identifier. An artificial enzyme identifier is created for each unique gene locus identifier. An auxotrophic mutant experiment is described as a query of $phenotypic\_effect/2$ whose outcome is a binary prediction. A $phenotypic\_effect/2$ query has a gene identifier and a list of reagents which is added to a list of pre-defined base growth medium compounds. These reagents are optional chemical substances that provide additional nutrients to facilitate metabolic reactions. Gene deletions are realised via $codes/2$ clauses where an enzyme corresponding to the deleted gene would be removed from the set of metabolic reactions to be enumerated.
@@ 7.2 Simulating a GEM using BMLP-IE
Output: A binary phenotypic effect classification.
@@
Another difference is that boolean matrices are created using the $reaction/2$ clauses for bottom-up evaluation. Algorithm 2 shows the application of BMLP-IE (Algorithm 1) for finding the transitive closure of synthesisable substrates. The two boolean matrices inputs to the BMLP-IE algorithm represent reactant and product substrates that participate in the reactions. The growth medium condition is encoded as the input vector to the BMLP-IE algorithm. This vector would be extended to represent substrates that are synthesisable by the organism in an auxotrophic mutant experiment. Specifying knockout genes excludes a subset of reactions. A vector computed during an intermediate step in the BMLP-IE algorithm denotes viable reactions given the set of knockout genes. A viable reaction is indicated by a binary bit 1 in this vector. Blocked reactions are excluded. In BMLP-IE, boolean operations compute the transition between metabolic states and return the closure set of synthesisable substrates. A binary phenotypic classification is made according to the inclusion of essential substrates in the closure set, indicating cell growth comparable to an unedited wild-type MG1655 E. coli strain. These essential substrates come from the biomass function of the wild-type in iML1515. The biomass function describes the necessary substrate composition for the growth of the organism.

### v3

@@ 3.1 Matrix algebra for datalog
We introduce datalog programs and concepts relating to our BMLP framework based on LP_book ; ILP_foundation . A variable is a character string beginning with an uppercase letter. A function, predicate or constant symbol is a character string starting with a lowercase letter. A term is either a variable, a constant symbol, or a function symbol attached to a tuple of terms. A ground term contains no variables. An atom is a predicate symbol applied to a tuple of terms e.g. $r(a,Y)$ where $r$ is a predicate symbol, $a$ is a constant symbol and $Y$ is a variable. An atom is ground when it contains only ground terms. A Horn clause is a disjunctive clause with at most one positive literal. A definite clause is a Horn clause with exactly one positive literal and a definite program is a set of definite clauses. A query is a user-posted Horn clause without any positive literal. A general logic program engine such as a Prolog interpreter can evaluate a query by carrying out a refutation proof of its negation via SLD resolutions (Selective Linear Definite clause resolution). A unit ground clause (fact) is a definite clause with no body literal and no variables. Datalog programs are definite programs without function symbols. A first-order definite clause has the form, $A_{0}\leftarrow A_{1},...,A_{n}$ where $A_{i}=r_{i}(x_{1},...,x_{k})$, $r_{i}$ is a predicate symbol, $x_{j}$ is a term. The arity of a function or predicate symbol is its number of arguments, e.g. $r_{i}/k$ indicates $r_{i}$ has arity $k$. The literal $A_{0}$ is called the head of the clause and the literals $A_{i}$ are the body.
@@
A recursive program has a body literal that appears in the head of a clause. Specifically, we focus on linear and immediately recursive datalog as simple recursive datalog programs. A linear and immediately recursive datalog program has a single recursive clause with the recursive predicate appearing in the body only once ioannidis_computation_1986 . This subset of recursive datalog programs can represent the relational structure of reaction pathways. This is further elaborated in Section 4. An example of a linear and immediately recursive datalog program involving one non-recursive predicate is:
@@
where unit ground clauses $r_{1}$ and a recursive clause $r_{2}$ can be represented as matrices R1, R2 such that $(\textbf{R}_{k})_{i,j}$ = 1 if $\mathcal{P}\models r_{k}(c_{i},c_{j})$ for constant symbols $c_{i},c_{j}$ and $(\textbf{R}_{k})_{i,j}$ = 0 otherwise. Addition “+” between two boolean matrices is defined as (A + B)i,j = Ai,j $\lor$ Bi,j. Multiplication “$\times$” is defined as $(\textbf{A}\times\textbf{B})_{i,j}$ = $\bigvee_{k=1}^{n}$ Ai,k $\land$ Bk,j and is abbreviated as AB.
@@ 3.2 Elementary nets
A marking $m^{\prime}$ is reachable from a marking $m$ if and only a finite step sequence exists. A place is reachable from $m$ if and only if $m^{\prime}$ is a reachable marking from $m$ and $m^{\prime}(p)=1$. A marking can be written as a row vector ${[}1,0,1,...{]}$ where each element indicates if the respective place has a token. A transition can also be represented by a row vector ${[}z_{1},z_{2},...,z_{k}{]}$ where $z_{i}$ reflects the assignment of tokens at the pre-set and post-set places. The consecutive firings of transition update markings and the global state of the system. This represents the dynamic changes in the distribution of tokenised resources.
@@ 4.1 Modelling a GEM with an OEN
We discuss the relationship between GEMs and OENs. Petri nets are commonly used formalism for representing complex biological systems such as genome-scale metabolic network models and diagnosing their dynamic behaviours sahu_advances_2021 . A genome-scale metabolic network model contains biochemical reactions of substances. Reactants are substances initially involved in a reaction and the products are the chemicals yielded by a reaction. A reaction is reversible when the conversion of reactants to products and products to reactants can take place simultaneously. Reversible reactions are commonly treated as two reactions in opposite directions. A reaction involves chemical substances $x_{i}$ and $y_{j}$:
@@
Given an auxotrophic mutant experiment, we model a GEM as an OEN to find the set of synthesisable substances by following the metabolic reaction pathways. This requires computing the necessary set of places for transitions to fire. Transitions need to be enumerated to examine if their necessary conditions can ever be met for them to fire. Every simulation of GEM involves a definition of the growth medium consisting of essential chemicals for cellular metabolism such as amino acids, inorganic salts and carbon sources. The definition of a growth medium serves as the basis of metabolic processes and gives an initial marking of OEN. Reachable places extending this marking are additional synthesisable substances.
@@ 4.2 Transforming an OEN into datalog
Each auxotrophic mutant experiment is treated as a datalog query to find all reachable places based on an initial marking. We evaluate queries of OEN reachability by transforming the OEN into a linear and immediately recursive datalog program. A datalog program consisting of clauses with arity two can be considered a directed graph where edges joining nodes represent the relation between terms. However, the multiplicity of the involved entities in a relation cannot be expressed this way. Instead, places and transitions are represented as constant and predicate symbols. Each argument of our transformed datalog program is a concatenation of constant symbols, e.g. $h2\_o2$. For our specific purpose, since all place nodes in the OEN are related to metabolites, every new concatenated constant symbol describes a set of metabolic substances.
This transformation with constant symbol concatenation generalises from a directed graph into a hypergraph berge_hypergraphs_1989 , a directed graph where an edge can connect to any number of nodes. Formally, a hypergraph $G=(V,E)$ is a tuple containing a set of nodes $V$ and a set of edges $E\subseteq 2^{V}\times 2^{V}$ between arbitrary subsets of $V$. Hypernodes are sets of nodes connected by such edges. The incidence graph of a hypergraph is bipartite in that for $v\in V$ and $e\in E$, $e$ is connected to $v$ if and only if $v$ and $e$ are nodes connected by an edge in the bipartite graph. Given an OEN, we consider a hypergraph with an incidence graph of the same bipartite structure. A hypernode represents a set of place nodes and is created for every transition’s input places from the OEN. Hypernodes are denoted by the new constant symbols that concatenate the member places’ names, e.g. a hypernode $\{H_{2},O_{2}\}$ is represented by the constant symbol $h2\_o2$. The number of hypernodes we create is fixed and limited by the number of transitions. This is because each reaction has a finite number of reactants and products. The dynamic behaviour of an OEN is specified via an initial marking. This mapping between an OEN and a hypergraph allows us to write a datalog program. A recursive clause is added which describes a path in the hypergraph.
@@
The right-hand-side figures are the OEN from from Figure 2 and the hypergraph represented by $\mathcal{P}_{1}$. The hypernode $\{c_{3},c_{4}\}$ is reachable from $\{c_{1},c_{2}\}$ via the transition $t_{1}$. The place $c_{5}$ is reachable in two steps via transitions $t_{1}$ and $t_{2}$. We can compute these by evaluating groundings of $pathway(m1,Y)$ in the Herbrand model $M(\mathcal{P}_{1})$ which gives us $pathway(m1,c2)$, $pathway(m1,c3\_c4)$ and $pathway(m1,c5)$.
@@ 5.1.1 Creating boolean matrices
We show boolean matrices created from $\mathcal{P}_{1}$ in Example 1. Individual substances are represented by constant symbols $c1$, $c2$ and $c3$. Sets of chemical substances are denoted by additional concatenated constant symbols, $c1\_c2$, $c3\_c4$. On the right-hand side, we create two boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$ from the reaction clauses in $\mathcal{P}_{1}$. Since the first reaction_1 clause also contains the constant symbols of the other reaction_1 clauses, to avoid redundancy, only one binary code is created for these clauses in $\textbf{R}_{1}$ and $\textbf{R}_{2}$.
@@
We show the mapping of a single row matrix from the token definition from $\mathcal{P}_{1}$ in Example 1. Individual substances are represented by constant symbols $c1$ and $c2$. The set of chemical substances is denoted by an additional concatenated constant symbol, $c1\_c2$.
@@ 5.2 Iterative extension (BMLP-IE)
Theorem 5.1 shows that BMLP-IE has a polynomial runtime cubic in the size of the boolean matrices. Evaluating the program $\mathcal{P}$ using SLD resolutions has the same time complexity and the runtime of BMLP-IE primarily benefits from bit-wise operations. For SLD resolutions, the size of the subset of the Herbrand base provable from a query $r(u,Y)$ is bounded by $O(n)$. Each member of this subset can be found via a depth-first search. Since there are $k$ unit ground clauses and each variable can be grounded by at most $n$ constants, this query requires at most $n$ search in a tree of $O(k\times n)$ nodes. For OEN reachability queries, it is important to note that the size of boolean matrices grows linearly with the number of hypernodes because they are both bounded by the number of transitions.
@@ 6.2 Expected cost of experiments
Experiment planning is often restricted by time and other resources. A rule of thumb for experiment design is to achieve optimal empirical outcomes within a finite budget. The problem of designing experiments to support or refute scientific hypotheses involves a series of binary decisions. This process is analogous to playing the game of “Guess Who” where players ask a sequence of “yes” or “no” questions whose answers binarily partition the candidate hypotheses. The outcome of an experiment $t$ splits hypotheses $H$ into consistent hypotheses $H_{t}$ and inconsistent hypotheses $\overline{H_{t}}$. The path of selection and outcomes of experiments can be represented as a binary decision tree. To construct an optimal tree, an active learner should seek to label instances consistent with up to half of the hypotheses mitchell_generalization_1982 .
@@ 6.3 Active learning sample complexity
The minimal reduction ratio of an actively selected instance can be computed before it is labelled. While in reality training examples might not be as discriminative, the optional selection strategy is to select instances with minimal reduction ratios as close as possible to $\frac{1}{2}$ with the ability to eliminate up to 50% of the hypothesis space. We describe the sample complexity advantage of active learning over random example selection by the following bound of an active version space learner’s sample complexity. A passive learner is considered a learner using random example selection since it does not have control over the training examples it uses.
@@ 7.1 Encoding a metabolic network
Our encoding of a metabolic network is based on the Robot Scientist King04:RobotScientist ; ase_progol . The Robot Scientist is implemented using the ILP system Progol progol and finds the synthesisable set of metabolites by traversing a metabolic network. It uses a metabolic network as a deductive database and answers a query of user-defined gene deletion and experimental conditions based on SLD resolutions. Its deductive database is written in Prolog syntax and uses Prolog list operations to manipulate constant symbols representing metabolic substances. Metabolic reactions are enumerated based on the set of substances currently synthesisable. The products of a metabolic reaction extend the list of synthesisable substances. Since the number of metabolic reactions is finite, the closure of members in this list needs to contain essential substances for the growth of an organism. In an auxotrophic mutant experiment outcome prediction, a gene function is removed. Some key metabolic reaction pathways might no longer be viable due to the removal, leading to insufficient substances for growth. In this case, the predicted experiment’s outcome is a lack of growth and a phenotypic effect.
In the Robot Scientist, the metabolic network is defined by $enzyme/6$ clauses. Each $enzyme/6$ clause contains an artificial identifier for an enzyme, gene locus identifiers, reaction identifiers, a reaction reversibility indicator and reactant and product identifiers. A gene locus corresponds to a physical site within a genome and is uniquely associated with a gene identifier. An artificial enzyme identifier is created for each unique gene locus identifier. An auxotrophic mutant experiment is described as a query of $phenotypic\_effect/2$ whose outcome is a binary prediction. A $phenotypic\_effect/2$ query has a gene identifier and a list of reagents which are optional chemical substances that provide additional nutrients to facilitate metabolic reactions. These reagents are added to a list of pre-defined base growth medium compounds. Gene deletions are realised via $codes/2$ clauses where an enzyme corresponding to the deleted gene would be removed from the set of metabolic reactions to be enumerated.
@@ 7.2 Simulating a GEM using BMLP-IE
Output: A binary phenotypic effect classification: 1 represents a phenotypic effect and 0 denotes no phenotypic effect.
@@
Another difference is that boolean matrices are created using the $reaction/2$ clauses for bottom-up evaluation. Algorithm 2 shows the application of BMLP-IE (Algorithm 1) for finding the transitive closure of synthesisable substances. The two boolean matrices inputs to the BMLP-IE algorithm represent reactants and products participating in the reactions. The growth medium condition is encoded as the input vector to the BMLP-IE algorithm. This vector would be extended to represent substances that are synthesisable by the organism in an auxotrophic mutant experiment. Specifying knockout genes excludes a subset of reactions. A vector computed during an intermediate step in the BMLP-IE algorithm denotes viable reactions given the set of knockout genes. A viable reaction is indicated by a binary bit 1 in this vector. Blocked reactions are excluded based on $codes/2$ and $enzyme/6$ definitions. In BMLP-IE, boolean operations compute the transition between metabolic states and return the closure set of synthesisable substances. A binary phenotypic classification is made according to the inclusion of essential substances in the closure set, indicating cell growth comparable to an unedited wild-type MG1655 E. coli strain. A binary classification of “1” corresponds to a phenotypic effect and “0” is no phenotypic effect. The essential substances come from the biomass function of the wild-type in iML1515 iML1515  [^2] We use the wild-type biomass function specified by iML1515 from http://bigg.ucsd.edu/ king_bigg_2016 . We use the qualitative version of this biomass function, meaning that we only care about the substances in the biomass, not their reaction coefficients.. The biomass function describes the necessary substance composition for the growth of the organism.